# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 5.2.7   The $\gamma^{*}p\to N(1680)\nicefrac{{5}}{{2}}^{+}$ transition
In the quark model, the $N(1680)\nicefrac{{5}}{{2}}^{+}$ is part of the $(\mathbf{56},2^{+})$ supermultiplet with spin $S=\nicefrac{{1}}{{2}}$ and orbital angular momentum $L=2$. It belongs to the same Regge trajectory as the $N(1520){\nicefrac{{3}}{{2}}^{-}}$, which is part of the $(\mathbf{70},1^{-})$ supermultiplet and also has three-quark spin $S=\nicefrac{{1}}{{2}}$ but one unit lower angular momentum $L=1$. The non-relativistic quark model with a harmonic oscillator potential predicts a very similar helicity structure for the ratios of $A_{1/2}$ and $A_{3/2}$ of the $N(1520)\nicefrac{{3}}{{2}}^{-}$ and the $N(1680)\nicefrac{{5}}{{2}}^{+}$ [949], indicating that the two resonance transitions should have a very similar helicity structure of the ratios (see, however, the comments in [950]):
$N(1520):\quad\frac{A_{1/2}}{A_{3/2}}=\frac{1}{\sqrt{3}}\left({\frac{{\mbox{\boldmath$Q$}}^{2}}{\alpha^{2}}}-1\right),\qquad N(1680):\quad\frac{A_{1/2}}{A_{3/2}}=-\frac{1}{2\sqrt{2}}\left(\frac{{\mbox{\boldmath$Q$}}^{2}}{\alpha^{2}}-2\right).$ (5.2)
Because of this similarity, we discuss the two resonance transitions together as we expect the same nature and a similar $Q^{2}$ dependence of the transition amplitudes for both states. The authors of Ref. [494, 651] consider the $N(1520)\nicefrac{{3}}{{2}}^{-}$ to be predominantly a composite meson-baryon resonance, while the $N(1680)\nicefrac{{5}}{{2}}^{+}$ is considered predominantly an elementary three-quark resonance. In this case, we should expect a different $Q^{2}$ behavior of their respective helicity amplitudes.
@@
Figure 5.15: Isovector and isoscalar part of $A_{1/2}(0)$ and $A_{3/2}(0)$ amplitudes for the $N(1675)\nicefrac{{5}}{{2}}^{-}$ [315].
Figure 5.16: Electroproduction transition amplitudes $A_{1/2}$, $A_{3/2}$ and $S_{1/2}$ for the $\gamma^{\ast}p\to N(1520)\nicefrac{{3}}{{2}}^{-}$ (top) and $\gamma^{\ast}p\to N(1680)\nicefrac{{5}}{{2}}^{+}$ (bottom) transitions. The curves are to guide the eyes. Source: Baryons reviews: $N$ and $\Delta$ Resonances in [951].
There are a number of nucleon and $\Delta$ excited states where electroproduction data for their transition amplitudes exist but are limited to narrow $Q^{2}$ ranges, e.g. $N(1650)\nicefrac{{1}}{{2}}^{-}$, $N(1710)\nicefrac{{1}}{{2}}^{+}$, $N(1880)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$, and $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$. Due to the limited coverage in $Q^{2}$ [18], these states are difficult to interpret in terms of the quark core and meson-baryon contributions to the excitation strength.
@@ 5.2.8   The $\gamma^{*}p\to N(1710)\nicefrac{{1}}{{2}}^{+}$ transition
The $N(1710)\nicefrac{{1}}{{2}}^{+}$ is the second excitation in the $J^{P}=\nicefrac{{1}}{{2}}^{+}$ nucleon channel, i.e., above the Roper resonance. This state is interpreted differently in the literature: In the AdS/QCD approach [702] it is the second radial excitation of the nucleon, while in Ref. [437] it was interpreted as dynamically generated resonance emerging from coupled-channel dynamics, jointly with a two-pole structure at the Roper mass. In Sec. 3.3 we discussed this resonance as a companion of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, which has one node but a different spatial distribution.
The insets in Fig. 5.17 shows quark-model calculations of these distributions for the $\Delta(1232)\nicefrac{{3}}{{2}}^{+},$ $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1710)\nicefrac{{1}}{{2}}^{+}$ as functions of the variables $\rho$ and $\lambda$ (in fm). The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ density distributions shows no significant structure. The $N(1440)\nicefrac{{1}}{{2}}^{+}$ density distribution shows a minimum as a function of $\rho^{2}+\lambda^{2}$, which is related to the zero crossing in the form factor $F_{2}$. The $N(1710)\nicefrac{{1}}{{2}}^{+}$ density distribution, on the other hand, shows a minimum as a function of $\rho^{2}-\lambda^{2}$. This is consistent with the data for $F_{2}$, which do not show any zero crossings as one might expect for a second radial excitation. Therefore, the $N(1710)\nicefrac{{1}}{{2}}^{+}$ seems to be incompatible with an interpretation as a $3s$ state, although higher statistics in the data would certainly be desirable. These observations also do not seem to support the interpretation in [437], where a parent state is supposed to generate three molecular states; if these states share the $qqq$ content of the bare state, they should also have similar transition form factors in electroproduction which is not the case. A study of a $qqq$ contribution to the three molecular states within the model of Ref. [437] could help resolve this issue.
@@
Figure 5.17: Transition form factor $F_{2}(Q^{2})$ for the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1710)\nicefrac{{1}}{{2}}^{+}$. The data are from Refs. [317, 311], the band is from [929] and the dashed curve from [382]. The quark-model spatial probability distributions from [668] are shown as insets.
@@ 5.3   What have we learned on the nature of the excited states from electroproduction?
The preceding discussions of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and $N(1535)\nicefrac{{1}}{{2}}^{-}$ electroproduction transition amplitudes show clear quark-core contributions that dominate the excitation strength at $Q^{2}>1.5-2$ GeV2. It is also interesting that the amplitudes are qualitatively compatible with the radial and orbital excitation structure expected from their assignments within the $SU(6)\times O(3)$ three-quark representation. This can be seen in Fig. 5.18, where the results from the single-quark transition model (SQTM) [952, 953, 954] for the transition amplitudes of the members of the $(SU(6),L^{P})$ = $(\mathbf{70},1^{-})$ supermultiplet are plotted. In this model, only single-quark transitions to the excited states are considered. After fitting three of the amplitudes to the data, the remaining 16 transverse helicity amplitudes for all states in $(\mathbf{70},1^{-})$ are determined, including transitions on neutrons. For resonances with sufficient reach in $Q^{2}$ the agreement is surprisingly good. For the neutrons, only data at the photon point are available, but they also show a surprisingly good correspondence. Note that no electroproduction measurements exist on neutron targets.
@@
Figure 5.18: Single-quark transition model (SQTM) amplitudes for the transitions from the ground-state $(\mathbf{56},0^{+})$ to the $(\mathbf{70},1^{-})$ supermultiplet for protons (left, blue) and neutrons (right, orange) [954, 15]. The markers are results of the data analysis. The colored bands for the top three amplitudes are parameterizations fitted to the data, whereas the bands in all other panels are SQTM predictions using those three as inputs. For the two $S$-wave states $N(1535)\nicefrac{{1}}{{2}}^{-}$ and $N(1650)\nicefrac{{1}}{{2}}^{-}$ a mixing angle of $31^{o}$ is used, and for the two $D$-wave states $N(1520)\nicefrac{{3}}{{2}}^{-}$ and $N(1700)\nicefrac{{3}}{{2}}^{-}$ the mixing angle of $6^{o}$ is applied.
@@
In general, all resonances we discussed in this section exhibit significant contributions in the lower $Q^{2}$ range that can be described in hadronic models as meson-baryon contributions. Here the transition $\gamma^{*}p\to N(1675)\nicefrac{{5}}{{2}}^{-}$ is particularly interesting as it is suppressed on proton targets via a selection rule, and we get a direct measurement of the meson-baryon contributions to the excitation of this resonance. These are found to be small at $Q^{2}>1.5$ GeV2 for $A_{3/2}$ and at $Q^{2}>3$ GeV2 for $A_{1/2}$, while both contributions, quark core and meson-baryon, are present in the case of neutron target $\gamma^{\ast}n\to N(1675)\nicefrac{{5}}{{2}}^{-}$ at the photon point.
@@
The knowledge of helicity amplitudes in a large $Q^{2}$ range also allows for the determination of transition charge densities on the light cone in transverse impact parameter space ($b_{x},b_{y}$[955, 378]. A comparison of the $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1535)\nicefrac{{1}}{{2}}^{-}$ light-front densities is shown in Figure 5.20. There are clear differences in the charge transition densities between the two states. The Roper has a softer positive core and a wider negative outer cloud than the $N(1535)\nicefrac{{1}}{{2}}^{-}$ and develops a larger shift in $b_{y}$ when the proton is polarized along the $b_{x}$ axis. These results are qualitatively consistent with the interpretation of a more extended spatial charge distribution from the widely distributed meson-baryon contributions to the Roper spatial wave function compared to the much narrower meson-baryon contributions to the $\gamma^{\ast}p\to N(1535)$ transition. The more narrowly shaped core contribution in the latter case also indicates that quark core contributions are strongly dominant over any meson-baryon contributions.
@@
Figure 5.20: (color online) Left panels: $N(1440)$, top: projection of charge densities on $b_{y}$, bottom: transition charge densities when the proton is spin polarized along $b_{x}$. Right panels: same for $N(1535)$. Note that the densities are scaled with $b^{2}$ to emphasize the outer wings. Color code:negative charge is blue, positive charge is red. Note that all scales are the same for ease of comparison [956]. Graphics credit: F.X. Girod.
@@
Despite the very significant progress made in recent years in further establishing the light-quark baryon spectrum and exploring the internal structure of excited states and the relationship to QCD [17, 287], much remains to be done. A large amount of precision data already collected needs to be included in the multi-channel analysis frameworks, and polarization data are still to be analyzed. There are approved proposals to study resonance excitations at much higher $Q^{2}$ and with higher precision at Jefferson Lab with CLAS12 [291, 957], which may begin to reveal the transition from a dressed quark core to a bare quark core at short distances. First attempts have been made on extending these coupled-channel analysis frameworks to the electroproduction data as well [448, 649].
@@
During the past decade, eight baryon states in the mass range from 1.85 to 2.15 GeV have been discovered, or evidence for the existence of states has been significantly strengthened. Some of these states are in the mass range and have $J^{PC}$ quantum numbers that could have significant contributions of gluonic components. Such “hybrid” states are in fact predicted in lattice QCD [810]. These states appear with the same quantum numbers as ordinary quark excitations and could possibly be isolated from ordinary states due to the $Q^{2}$ dependence of their helicity amplitudes [958], which is expected to be quite different from ordinary 3-quark excitation. The study of hybrid baryon excitations then requires new electroproduction data, especially at low $Q^{2}$ [959] with different final states and masses above 2 GeV.
@@
A novel way of studying the structure of the proton has been experimentally pioneered in recent years and resulted in the first determination of the mechanical properties of the proton, its internal distribution of pressure and shear stress [960]. This remarkable result is based on the similarity of the graviton coupling, a spin-2 interaction, with the coupling of two spin-1 photons in the exclusive process of deeply virtual Compton scattering (DVCS). For a colloquial introduction in this topic, see [961]. If energy is transferred to the ground-state proton, a highly energetic photon may be released while the proton remains in its ground state, in which case the process probes the generalized parton distributions (GPDs) of the nucleon. If an intermediate nucleon excitation is formed, which then decays into a nucleon and a pion or some other final state in addition to the photon, the process probes the $NN^{*}$ transition GPDs from the nucleon to the intermediate resonance. Recently, the process $ep\to e\pi^{-}\Delta^{++}$, which is sensitive to the $N\to\Delta$ transition GPDs, has been measured [962], see [963] for a review. While this research is still in its very early stage, much more progress can be expected in the coming years. It is worth mentioning that the DVCS process in Fig. 5.20 is a hard, deep inelastic process, where the final-state photon has energies $E_{\gamma}\geq 2-3$ GeV, and the reaction provides access to the transition GPDs of the resonance’s quark core.
@@
Where does one go from here on the theory side? Despite the substantial progress discussed so far, the theoretical description of baryon resonances remains at an early stage. As we have seen, current modelling allows one to separate the physics into a three-quark core at high $Q^{2}$ and meson-baryon interactions at low $Q^{2}$, with various approaches combining these components in a suitable framework. In this section we focused on recent results for nucleon resonance transition amplitudes and their interpretation within different approaches, including light-front relativistic quark models with momentum-dependent constituent quark masses, light-cone sum rules [964], and functional methods based on Dyson-Schwinger equations [965, 692]; for further examples see [18] and references therein. Typically, such calculations describe the transition form factors at $Q^{2}\gtrsim 2$ GeV2, while at lower $Q^{2}$ values hadronic degrees of freedom must be included. These may account for significant parts of the total resonance excitation strength.

### v4

@@ 5.2.7  The $\gamma^{*}p\to N(1680)\nicefrac{{5}}{{2}}^{+}$ transition
In the quark model, the $N(1680)\nicefrac{{5}}{{2}}^{+}$ is part of the $(\mathbf{56},2^{+})$ supermultiplet with spin $S=\nicefrac{{1}}{{2}}$ and orbital angular momentum $L=2$. It belongs to the same Regge trajectory as the $N(1520){\nicefrac{{3}}{{2}}^{-}}$, which is part of the $(\mathbf{70},1^{-})$ supermultiplet and also has three-quark spin $S=\nicefrac{{1}}{{2}}$ but one unit lower angular momentum $L=1$. The non-relativistic quark model with a harmonic oscillator potential predicts a very similar helicity structure for the ratios of $A_{1/2}$ and $A_{3/2}$ of the $N(1520)\nicefrac{{3}}{{2}}^{-}$ and the $N(1680)\nicefrac{{5}}{{2}}^{+}$ Copley:1969qn , indicating that the two resonance transitions should have a very similar helicity structure of the ratios (see, however, the comments in Ono:1976mx ):
$\hskip-17.07164ptN(1520):\quad\frac{A_{1/2}}{A_{3/2}}=\frac{1}{\sqrt{3}}\left({\frac{{\mbox{$Q$}}^{2}}{\alpha^{2}}}-1\right),\qquad N(1680):\quad\frac{A_{1/2}}{A_{3/2}}=-\frac{1}{2\sqrt{2}}\left(\frac{{\mbox{$Q$}}^{2}}{\alpha^{2}}-2\right).$ (5.2)
Because of this similarity, we discuss the two resonance transitions together as we expect the same nature and a similar $Q^{2}$ dependence of the transition amplitudes for both states. The authors of Ref. Wang:2024jns , Wang:2023snv consider the $N(1520)\nicefrac{{3}}{{2}}^{-}$ to be predominantly a composite meson-baryon resonance, while the $N(1680)\nicefrac{{5}}{{2}}^{+}$ is considered predominantly an elementary three-quark resonance. In this case, we should expect a different $Q^{2}$ behavior of their respective helicity amplitudes.
@@
Figure 5.15: Isovector and isoscalar part of $A_{1/2}(0)$ and $A_{3/2}(0)$ amplitudes for the $N(1675)\nicefrac{{5}}{{2}}^{-}$ Aznauryan:2014xea .
Figure 5.16: Electroproduction transition amplitudes $A_{1/2}$, $A_{3/2}$ and $S_{1/2}$ for the $\gamma^{\ast}p\to N(1520)\nicefrac{{3}}{{2}}^{-}$ (top) and $\gamma^{\ast}p\to N(1680)\nicefrac{{5}}{{2}}^{+}$ (bottom) transitions. The curves are to guide the eyes. Source: Baryons reviews: NN and $\Delta$ Resonances in Workman:2022ynf .
There are a number of nucleon and $\Delta$ excited states where electroproduction data for their transition amplitudes exist but are limited to narrow $Q^{2}$ ranges, e.g. $N(1650)\nicefrac{{1}}{{2}}^{-}$, $N(1710)\nicefrac{{1}}{{2}}^{+}$, $N(1880)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$, and $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$. Due to the limited coverage in $Q^{2}$ Ramalho:2023hqd , these states are difficult to interpret in terms of the quark core and meson-baryon contributions to the excitation strength.
@@ 5.2.8  The $\gamma^{*}p\to N(1710)\nicefrac{{1}}{{2}}^{+}$ transition
The $N(1710)\nicefrac{{1}}{{2}}^{+}$ is the second excitation in the $J^{P}=\nicefrac{{1}}{{2}}^{+}$ nucleon channel, i.e., above the Roper resonance. This state is interpreted differently in the literature: In the AdS/QCD approach Brodsky:2014yha it is the second radial excitation of the nucleon, while in Ref. Suzuki:2009nj it was interpreted as dynamically generated resonance emerging from coupled-channel dynamics, jointly with a two-pole structure at the Roper mass. In Sec. 3.3 we discussed this resonance as a companion of the $N(1440)\nicefrac{{1}}{{2}}^{+}$, which has one node but a different spatial distribution.
The insets in Fig. 5.17 shows quark-model calculations of these distributions for the $\Delta(1232)\nicefrac{{3}}{{2}}^{+},$ $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1710)\nicefrac{{1}}{{2}}^{+}$ as functions of the variables $\rho$ and $\lambda$ (in fm). The $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ density distributions shows no significant structure. The $N(1440)\nicefrac{{1}}{{2}}^{+}$ density distribution shows a minimum as a function of $\rho^{2}+\lambda^{2}$, which is related to the zero crossing in the form factor $F_{2}$. The $N(1710)\nicefrac{{1}}{{2}}^{+}$ density distribution, on the other hand, shows a minimum as a function of $\rho^{2}-\lambda^{2}$. This is consistent with the data for $F_{2}$, which do not show any zero crossings as one might expect for a second radial excitation. Therefore, the $N(1710)\nicefrac{{1}}{{2}}^{+}$ seems to be incompatible with an interpretation as a $3s$ state, although higher statistics in the data would certainly be desirable. These observations also do not seem to support the interpretation in Suzuki:2009nj , where a parent state is supposed to generate three molecular states; if these states share the qqqqqq content of the bare state, they should also have similar transition form factors in electroproduction which is not the case. A study of a qqqqqq contribution to the three molecular states within the model of Ref. Suzuki:2009nj could help resolve this issue.
@@
Figure 5.17: Transition form factor $F_{2}(Q^{2})$ for the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1710)\nicefrac{{1}}{{2}}^{+}$. The data are from Refs. CLAS:2012wxw , CLAS:2014fml , the band is from Eichmann:2018ytt and the dashed curve from Drechsel:2007if . The quark-model spatial probability distributions from Melde:2008yr are shown as insets.
@@ 5.3  What have we learned on the nature of the excited states from electroproduction?
The preceding discussions of the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$, $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and $N(1535)\nicefrac{{1}}{{2}}^{-}$ electroproduction transition amplitudes show clear evidence of quark-core contributions, particularly at $Q^{2}>2$ GeV2 where they nearly coincide within uncertainties with the quark-based approaches of the LFRQM, DSE and LCSR calculations. Furthermore, for the $N(1675)\nicefrac{{5}}{{2}}^{-}$ state we find a quantitative estimate of the non-quark contribution to the proton transition amplitudes $A_{1/2}(Q^{2})$ and $A_{3/2}(Q^{2})$ owing to the suppression of the quark contribution. For the neutron an estimate at the real photon point leads to results that are similar to the proton albeit with opposite sign. It indicates the isovector character of these contributions.
@@
It is also interesting that the amplitudes are qualitatively compatible with the radial and orbital excitation structure expected from their assignments within the $SU(6)\times O(3)$ three-quark representation. This can be seen in Fig. 5.18, where the results from the algebraic single-quark transition model (SQTM) Hey:1974qe , Cottingham:1978za , Burkert:2002zz for the transition amplitudes of the members of the $(SU(6),L^{P})$ = $(\mathbf{70},1^{-})$ supermultiplet are plotted. In this model, only single-quark transitions to the excited states are considered. After fitting three of the amplitudes to the data, the remaining 16 transverse helicity amplitudes for all states in $(\mathbf{70},1^{-})$ are determined, including transitions on neutrons. For resonances with sufficient reach in $Q^{2}$ the agreement is surprisingly good. For the neutrons, only data at the photon point are available, but they also show a surprisingly good correspondence.
@@
Figure 5.18: Single-quark transition model (SQTM) amplitudes for the transitions from the ground-state $(\mathbf{56},0^{+})$ to the $(\mathbf{70},1^{-})$ supermultiplet for protons (left, blue) and neutrons (right, orange) Burkert:2002zz , Aznauryan:2011qj . The markers are results of the data analysis. The colored bands for the top three amplitudes are parameterizations fitted to the data, whereas the bands in all other panels are SQTM predictions using those three as inputs. For the two $S$-wave states $N(1535)\nicefrac{{1}}{{2}}^{-}$ and $N(1650)\nicefrac{{1}}{{2}}^{-}$ a mixing angle of $31^{o}$ is used, and for the two $D$-wave states $N(1520)\nicefrac{{3}}{{2}}^{-}$ and $N(1700)\nicefrac{{3}}{{2}}^{-}$ the mixing angle of $6^{o}$ is applied. Note that these graphs do not include more recent data covering higher $Q^{2}$ results.
@@
In general, all resonances we discussed in this section exhibit significant contributions in the lower $Q^{2}$ range that may be described in hadronic models as meson-baryon contributions. Here the transition $\gamma^{*}p\to N(1675)\nicefrac{{5}}{{2}}^{-}$ is particularly interesting as it is suppressed on proton targets via a selection rule, and we get a direct measurement of the meson-baryon contributions to the excitation of this resonance. These are found to be small at $Q^{2}>1.5$ GeV2 for $A_{3/2}$ and at $Q^{2}>3$ GeV2 for $A_{1/2}$, while both contributions, quark core and meson-baryon, are present in the case of neutron target $\gamma^{\ast}n\to N(1675)\nicefrac{{5}}{{2}}^{-}$ at the photon point.
The experimental data used in Fig. 5.8 and Figs. 5.105.16 were recently also analyzed by the Jülich-Bonn-Washington (JBW) group  Wang:2024byt . The group employs a dynamical coupled channel approach. The results are presented as the $Q^{2}$ dependence of the imaginary and real part of the helicity transition amplitudes $A_{1/2},~A_{3/2}$ and $S_{1/2}$. The real parts compare generally well with the results of the data analysis summarized in  Aznauryan:2011qj with the exception of the $A_{1/2}$ amplitude of $N(1440)1/2^{+}$ at very small $Q^{2}$, where the zero crossing is observed at different $Q^{2}$. This is an interesting discrepancy and should be studied further; it does, however, not affect the conclusions we draw in this article.
The Argonne-JLab-Osaka group has determined some of the resonance transition amplitudes employing a dynamical coupled channel approach Suzuki:2010yn . They extract transition amplitudes for two Roper-like states at pole masses just 8 MeV apart from each other. If one adds the real parts of the two amplitudes, they compare qualitatively well with the results we discuss in this article. Similar observations apply for the $N(1520)\nicefrac{{3}}{{2}}^{-}$, where the results compare quantitatively well in the limited $Q^{2}$ range that was analyzed by the group.
@@
The knowledge of helicity amplitudes in a large $Q^{2}$ range also allows for the determination of transition charge densities on the light cone in transverse impact parameter space ($b_{x},b_{y}$Carlson:2007xd , Tiator:2008kd . A comparison of the $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $N(1535)\nicefrac{{1}}{{2}}^{-}$ light-front densities is shown in Figure 5.20. There are clear differences in the charge transition densities between the two states. The Roper has a softer positive core and a wider negative outer cloud than the $N(1535)\nicefrac{{1}}{{2}}^{-}$ and develops a larger shift in $b_{y}$ when the proton is polarized along the $b_{x}$ axis. These results are qualitatively consistent with the interpretation of a more extended spatial charge distribution from the widely distributed meson-baryon contributions to the Roper spatial wave function compared to the much narrower meson-baryon contributions to the $\gamma^{\ast}p\to N(1535)$ transition. The more narrowly shaped core contribution in the latter case also indicates that quark core contributions are strongly dominant over any meson-baryon contributions.
@@
Figure 5.20: (color online) Left panels: $N(1440)$, top: projection of charge densities on $b_{y}$, bottom: transition charge densities when the proton is spin polarized along $b_{x}$. Right panels: same for $N(1535)$. Note that the densities are scaled with $b^{2}$ to emphasize the outer wings. Color code:negative charge is blue, positive charge is red. Note that all scales are the same for ease of comparison Burkert:2018oyl . Graphics credit: F.X. Girod.
@@
Despite the very significant progress made in recent years in further establishing the light-quark baryon spectrum and exploring the internal structure of excited states and the relationship to QCD Carman:2020qmb , Proceedings:2020fyd , much remains to be done. A large amount of precision data already collected needs to be included in the multi-channel analysis frameworks, and polarization data are still to be analyzed. There are approved proposals to study resonance excitations at much higher $Q^{2}$ and with higher precision at Jefferson Lab with CLAS12 CLAS:2022kta , Burkert:2018nvj , which may begin to reveal the transition from a dressed quark core to a bare quark core at short distances. First attempts have been made on extending these coupled-channel analysis frameworks to the electroproduction data as well Mai:2023cbp , Wang:2024byt .
@@
During the past decade, eight baryon states in the mass range from 1.85 to 2.15 GeV have been discovered, or evidence for the existence of states has been significantly strengthened. Some of these states are in the mass range and have $J^{PC}$ quantum numbers that could have significant contributions of gluonic components. Such “hybrid” states are in fact predicted in lattice QCD Dudek:2012ag . These states appear with the same quantum numbers as ordinary quark excitations and could possibly be isolated from ordinary states due to the $Q^{2}$ dependence of their helicity amplitudes Li:1991yba , which is expected to be quite different from ordinary 3-quark excitation. The study of hybrid baryon excitations then requires new electroproduction data, especially at low $Q^{2}$ Lanza:2021ayj with different final states and masses above 2 GeV.
@@
A novel way of studying the structure of the proton has been experimentally pioneered in recent years and resulted in the first determination of the mechanical properties of the proton, its internal distribution of pressure and shear stress Burkert:2018bqq . This remarkable result is based on the similarity of the graviton coupling, a spin-2 interaction, with the coupling of two spin-1 photons in the exclusive process of deeply virtual Compton scattering (DVCS). For a colloquial introduction in this topic, see  Burkert:2023wzr . If energy is transferred to the ground-state proton, a highly energetic photon may be released while the proton remains in its ground state, in which case the process probes the generalized parton distributions (GPDs) of the nucleon. If an intermediate nucleon excitation is formed, which then decays into a nucleon and a pion or some other final state in addition to the photon, the process probes the $NN^{*}$ transition GPDs from the nucleon to the intermediate resonance. Recently, the process $ep\to e\pi^{-}\Delta^{++}$, which is sensitive to the $N\to\Delta$ transition GPDs, has been measured CLAS:2023akb , see Diehl:2024bmd for a review. While this research is still in its very early stage, much more progress can be expected in the coming years. It is worth mentioning that the DVCS process in Fig. 5.20 is a hard, deep inelastic process, where the final-state photon has energies $E_{\gamma}\geq 2-3$ GeV, and the reaction provides access to the transition GPDs of the resonance’s quark core.
@@
Where does one go from here on the theory side? Despite the substantial progress discussed so far, the theoretical description of baryon resonances remains at an early stage. As we have seen, current modelling allows one to separate the physics into a three-quark core at high $Q^{2}$ and meson-baryon interactions at low $Q^{2}$, with various approaches combining these components in a suitable framework. In this section we focused on recent results for nucleon resonance transition amplitudes and their interpretation within different approaches, including light-front relativistic quark models with momentum-dependent constituent quark masses, light-cone sum rules Braun:2009jy , and functional methods based on Dyson-Schwinger equations Roberts:2007ji , Eichmann:2016yit ; for further examples see Ramalho:2023hqd and references therein. Typically, such calculations describe the transition form factors at $Q^{2}\gtrsim 2$ GeV2, while at lower $Q^{2}$ values hadronic degrees of freedom must be included. These may account for significant parts of the total resonance excitation strength.